# Length-Controlled Margin-Based Preference Optimization without Reference Model

### v1

@@ Abstract
Direct Preference Optimization (DPO) is a widely adopted offline algorithm for preference-based reinforcement learning from human feedback (RLHF), designed to improve training simplicity and stability by redefining reward functions. However, DPO is hindered by several limitations, including length bias, memory inefficiency, and probability degradation. To address these challenges, we propose Length-Controlled Margin-Based Preference Optimization (LMPO), a more efficient and robust alternative. LMPO introduces a uniform reference model as an upper bound for the DPO loss, enabling a more accurate approximation of the original optimization objective. Additionally, an average log-probability optimization strategy is employed to minimize discrepancies between training and inference phases. A key innovation of LMPO lies in its Length-Controlled Margin-Based loss function, integrated within the Bradley-Terry framework. This loss function regulates response length while simultaneously widening the margin between preferred and rejected outputs. By doing so, it mitigates probability degradation for both accepted and discarded responses, addressing a significant limitation of existing methods. We evaluate LMPO against state-of-the-art preference optimization techniques on two open-ended large language models, Mistral and LLaMA3, across six conditional benchmarks. Our experimental results demonstrate that LMPO effectively controls response length, reduces probability degradation, and outperforms existing approaches. The code is available at https://github.com/gengxuli/LMPO.
@@ 1 Introduction
Human feedback is essential for aligning large language models (LLMs) with human values and objectives (Jiang et al., 2024; Chang et al., 2024), ensuring that these models act in ways that are helpful, reliable, and safe. A common strategy for achieving this alignment is reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022), which fine-tunes language models using human evaluations. While RLHF has shown substantial success (Schulman et al., 2017), it also introduces notable challenges in optimization due to its multi-step design. This process first involves training a reward model to evaluate outputs based on human preferences, and then optimizing a policy model to maximize the assigned rewards. The complexity of these sequential steps often complicates the implementation and reduces efficiency (Chaudhari et al., 2024).
@@
In response to these challenges, researchers have started exploring simpler alternatives that avoid the intricate, multi-stage nature of RLHF. One promising method is Direct Preference Optimization (DPO) (Rafailov et al., 2024), which streamlines the process by reformulating the reward function. This approach enables direct learning of a policy model from preference data, eliminating the need for a separate reward model. As a result, DPO offers greater stability and is more practical to implement.
DPO estimates implicit rewards using the log-probability ratio between a policy model’s response and that of a supervised fine-tuned (SFT) model, enabling preference learning without an explicit reward function. However, this implicit reward may misalign with the log-probability metric during inference. Moreover, DPO’s reliance on both policy and SFT models significantly increases GPU usage, especially for LLMs. The DPO loss, derived from the Bradley-Terry model, can create training imbalances, as it does not ensure an increase in the probability of positive samples—potentially reducing both positive and negative probability simultaneously. Unlike IPO   (Azar et al., 2024), which constrains probability variation but weakens response distinction, DPO also exhibits length bias, favoring longer responses due to preference label distribution inconsistencies (Lu et al., 2024). This issue, common in multi-stage RLHF methods, allows models to exploit verbosity for higher rewards without improving output quality, often generating responses nearly twice as long as labeled data.
@@
  • Competitive performance: Despite being reference-free, our method demonstrates competitive performance when compared to DPO and its variants (Hong et al., 2024a; Ethayarajh et al., 2024). This performance advantage is consistent across a variety of training setups and comprehensive instruction-following benchmarks, including AlpacaEval 2 (Li et al., 2023) and Arena-Hard v0.1 (Li et al., 2024).
@@ 2 Related Work
Reinforcement learning with human feedback (RLHF) often utilizes the Bradley-Terry model (Bradley and Terry, 1952) to estimate the probability of success in pairwise comparisons between two independently evaluated instances. Additionally, a reward model is trained to assign scores to these instances. Reinforcement learning algorithms, such as proximal policy optimization (PPO) (Schulman et al., 2017), are used to train models to maximize the reward model’s score for the selected response, ultimately enabling LLMs to align with human preferences (Stiennon et al., 2020; Ziegler et al., 2019). A notable example is InstructGPT (Ouyang et al., 2022), which showcased the scalability and adaptability of RLHF in training instruction-following language models. Alternative approaches, such as reinforcement learning with language model feedback (RLAIF (Lee et al., 2023)), may also serve as feasible substitutes for human feedback (Bai et al., 2022; Sun et al., 2023). Nevertheless, RLHF encounters challenges, including the need for extensive hyperparameter tuning due to the instability of PPO (Rafailov et al., 2024) and the sensitivity of the reward models (Wang et al., 2024). Consequently, there is a pressing demand for more stable preference alignment algorithms.
@@
Several techniques for preference alignment reduce the reliance on reinforcement learning. Direct Policy Optimization (DPO) (Rafailov et al., 2024) is a method that integrates reward modeling with preference learning. And Identity Preference Optimization (IPO) (Azar et al., 2024) is introduced to mitigate potential overfitting issues in DPO. In contrast to RLHF and DPO, an alternative approach called Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) is proposed, which does not require pairwise preference datasets. Additionally, Preference Ranking Optimization (PRO) (Song et al., 2024) introduces the incorporation of the softmax values from the reference response set into the negative log-probability (NLL) loss, allowing for a unified approach to supervised fine-tuning and preference alignment.
@@
Due to the reliance of DPO and DPO-like methods on both the policy model and the SFT model during the alignment process, they impose greater demands on GPU resources. Several techniques have been developed to alleviate this GPU requirement by eliminating the need for a reference model. CPO (Xu et al., 2024) demonstrates that the ideal loss function without a reference model can serve as the upper bound of the DPO loss, with the SFT loss acting as a replacement for the KL divergence. ORPO (Hong et al., 2024a) models the optimal reward as a log-odds function, removing the need for an additional fixed reference model. MaPO (Hong et al., 2024b) builds on the ORPO approach by introducing a margin-aware term for aligning diffusion models without a reference model. SimPO (Meng et al., 2024) adopts a similar reference-free preference learning framework as CPO but with improved stability due to its specific length normalization and target reward margin, leading to superior performance in various benchmarks.
@@ 3.1 Direct Preference Optimization (DPO)
We derive our method by first examining DPO (Rafailov et al., 2024), which provides a more straightforward optimization goal within the framework of RLHF (Ziegler et al., 2019; Stiennon et al., 2020). DPO operates on a dataset of source sentences, $x$, paired with both preferred translations, $y_{w}$, and less preferred ones, $y_{l}$. This dataset, containing comparison examples, is denoted as $\mathcal{D}=\left\{x^{(i)},y_{w}^{(i)},y_{l}^{(i)}\right\}_{i=1}^{N}$. The loss function for DPO is formulated as a maximum likelihood estimation for a policy model parameterized by $\pi_{\theta}$:
@@
where $\pi_{\text{ref}}$ refers to a SFT model, $\sigma$ represents the sigmoid function, and $\beta$ is a scaling hyperparameter. The formulation of the DPO loss is based on a reparameterization of the true reward signal and the corresponding optimal policy, borrowing from the PPO framework (Schulman et al., 2017). This loss allows DPO to be trained in a supervised fine-tuning manner, as it makes exclusive use of labeled preference data without requiring any interaction between the agent and its environment which is a shortcoming for PPO.
3.2 Revisiting Bradley-Terry Model
DPO in Section 3.1 uses a statistical model commonly used for sporting events called Bradley-Terry. The Bradley-Terry model stipulates that the human preference distribution $p^{*}$ can be written as:
@@ 3.2 Revisiting Bradley-Terry Model
The BT model used in DPO is the original form. There are some variants that make some improvements on the BT model. Rao-Kupper model (Rao and Kupper, 1967) considers model human preference with ties: $p^{*}(y_{w}=y_{l}\mid x)$, which means two responses $(y_{w},y_{l})$ are considered equal with respect to the prompt $x$.
So in order to better distinguish the two responses, we define the loss response as a home-filed team in the BT model. And we may incorporate a home-court advantage by including an intercept term $h$:
@@
For DPO, $d(x,y_{w},y_{l})$ means the term in function $\sigma$, which is outlined in Section 3.1. DPO mitigates several issues inherent in conventional RLHF techniques and has found widespread application in modern models, including Meta’s recently released Llama 3.1 model (Dubey et al., 2024). Despite these advantages, DPO presents notable drawbacks when compared to standard supervised fine-tuning. One major limitation is its inefficiency in memory usage, as it requires doubling the memory to accommodate both the trained policy and the reference policy concurrently. Additionally, DPO suffers from reduced computational efficiency, as the model must be executed separately for each policy, effectively doubling the processing time. So it is of vital importance to investigate a reference model-free RLHF method.
A recent method called CPO(Xu et al., 2024) has proved that when $\pi_{\text{ref}}$ is defined as $\pi_{w}$, an ideal policy that precisely aligns with the true data distribution of preferred data, the DPO loss $\mathcal{L}(\pi_{\theta};\pi_{w})+C$ is upper bounded by $\mathcal{L}(\pi_{\theta};U)$, where $C$ is a constant. So following this proof, we use a uniform reference model to approximate $d(x,y_{w},y_{l})$:
@@
Next, in DPO, the implicit reward is formulated using the log ratio of the probability of a response between the current policy model and the SFT model. However, this reward formulation is not directly aligned with the metric used to guide generation, which is approximately the average log probability of a response generated by the policy model. So there is an assumption that this discrepancy between training and inference phases may lead to bad performance. In order to eliminate this discrepancy, we replace the log probability with the average log probability in Eq. 4:
@@ 3.3 Length-Controlled Margin-Based Loss
Here, $m(y_{w},y_{l},x)$ represents a margin that quantifies the preference strength between the winning response $y_{w}$ and the losing response $y_{l}$ for a given input $x$, while $\lambda$ is a scaling factor. The function $r^{*}(x,y)$ provides the reward score for response $y$ conditioned on input prompt $x$. By including this margin, the model is better able to differentiate reward scores, especially when the quality gap between responses is substantial.
Recent approaches have adopted this formulation to enhance model performance. For example, the reward models in Llama-2-Chat (Touvron et al., 2023) and UltraRM (Cui et al., 2023) use discrete preference scores as margin terms. SimPO (Meng et al., 2024) employs a fixed margin to guarantee that the reward for the preferred response always exceeds that of the less favored one. Despite these advances, issues such as length bias persist.
In response to this issue, we introduce the Length-Controlled Margin-Based Loss, which is designed to address several key limitations. First, it explicitly controls the length of generated responses, thereby mitigating the bias towards longer outputs often seen in LLMs. Additionally, the loss function regulates the probability decrease for both selected and rejected responses, further ensuring that the model can more clearly distinguish between correct and incorrect responses. Importantly, this framework also aims to increase the margin between the probabilities of chosen and rejected responses, thus amplifying the model’s capacity to discriminate between high- and low-quality responses. The full formulation of the Length-Controlled Margin-Based Loss is presented below.
@@
Objective. Finally, we obtain the LMPO finall loss function by incorporating the above considerations:
@@
In summary, LMPO employs an implicit reward formulation that directly aligns with the generation metric, eliminating the need for a reference model. Next, it introduces a margin term $m(\mathbf{x},\mathbf{y}^{w},\mathbf{y}^{l})$ with two kinds of normalization methods to help separate the winning and losing responses, alleviate length bias and wining response probability decrement problems.
@@ 4.1 Experimental Setup
Models and training settings. We perform preference optimization with two families of models, Llama3-8B(AI@Meta, 2024) and Mistral-7B(Jiang et al., 2023) under two setups: Base and Instruct.
For the Base experimental setup, following SimPO, we utilize pre-trained models (alignment-handbook/zephyr-7b-sft-full(Tunstall et al., 2023) and (princeton-nlp/Llama-3-Base-8B-SFT) as SFT models. These SFT models are then used as the foundation for preference optimization on the UltraFeedback dataset (Cui et al., 2023), which collects feedback ($y_{w}$ and $y_{l}$) from LLMs of different quality levels.
For the Instruct experimental setup, we utilize pre-trained instruction-tuned models (mistralai/Mistral-7B-Instruct-v0.2) and (meta-llama/Meta-Llama-3-8B-Instruct) as SFT models. For a fair comparison, we use the same training data as SimPO: (princeton-nlp/llama3-ultrafeedback) and (https://huggingface.co/datasets/princeton-nlp/mistral-instruct-ultrafeedback) for Llama3-8B and Mistral-7B, respectively.
These configurations embody the latest advancements, securing our models a place among the top contenders on various leaderboards.
@@
Additionally, we evaluate the models on six downstream tasks in the Huggingface Open Leaderboard V1, following SimPO (Meng et al., 2024). These downstream tasks include the AI2 Reasoning Challenge (25-shot) (Clark et al., 2018), HellaSwag (10-shot) (Zellers et al., 2019), MMLU (5-shot) (Hendrycks et al., 2020), TruthfulQA (0-shot) (Lin et al., 2021), Winogrande (5-shot) (Sakaguchi et al., 2021), and GSM8K (5-shot) (Cobbe et al., 2021). We report the match accuracy for these conditional benchmarks. Additional details are provided in Appendix A.
Baselines We perform a comparative analysis of our method against several state-of-the-art offline preference optimization techniques, including DPO (Rafailov et al., 2024), IPO (Azar et al., 2024), CPO (Xu et al., 2024), KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024a), R-DPO (Park et al., 2024), and SimPO (Meng et al., 2024). For SimPO, we use the model provided for the Llama3-8B family and replicate the SimPO methodology for the Mistral-7B family in our environment. For the other methods, we report the results provided by SimPO. We also tune the hyperparameters for SimPO and report the best performance achieved.

### v2

@@ Abstract
Direct Preference Optimization (DPO) is a widely adopted offline algorithm for preference-based reinforcement learning from human feedback (RLHF), designed to improve training simplicity and stability by redefining reward functions. However, DPO is hindered by several limitations, including length bias, memory inefficiency, and probability degradation. To address these challenges, we propose Length-Controlled Margin-Based Preference Optimization (LMPO), a more efficient and robust alternative. LMPO introduces a uniform reference model as an upper bound for the DPO loss, enabling a more accurate approximation of the original optimization objective. Additionally, an average log-probability optimization strategy is employed to minimize discrepancies between training and inference phases. A key innovation of LMPO lies in its Length-Controlled Margin-Based loss function, integrated within the Bradley-Terry framework. This loss function regulates response length while simultaneously widening the margin between preferred and rejected outputs. By doing so, it mitigates probability degradation for both accepted and discarded responses, addressing a significant limitation of existing methods. We evaluate LMPO against state-of-the-art preference optimization techniques on two open-ended large language models, Mistral and LLaMA3, across ten conditional benchmarks and two open-ended benchmarks. Our experimental results demonstrate that LMPO effectively controls response length, reduces probability degradation, and outperforms existing approaches. The code is available at https://github.com/gengxuli/LMPO.
@@ 1 Introduction
Human feedback is essential for aligning large language models (LLMs) with human values and objectives Jiang et al. (2024); Chang et al. (2024), ensuring that these models act in ways that are helpful, reliable, and safe. A common strategy for achieving this alignment is reinforcement learning from human feedback (RLHF) Ziegler et al. (2019); Stiennon et al. (2020); Ouyang et al. (2022), which fine-tunes language models using human evaluations. While RLHF has shown substantial success Schulman et al. (2017), it also introduces notable challenges in optimization due to its multi-step design. This process first involves training a reward model to evaluate outputs based on human preferences, and then optimizing a policy model to maximize the assigned rewards. The complexity of these sequential steps often complicates the implementation and reduces efficiency Chaudhari et al. (2024).
In response to these challenges, researchers have started exploring simpler alternatives that avoid the intricate, multi-stage nature of RLHF. One promising method is Direct Preference Optimization (DPO) Rafailov et al. (2024), which streamlines the process by reformulating the reward function. This approach enables direct learning of a policy model from preference data, eliminating the need for a separate reward model. As a result, DPO offers greater stability and is more practical to implement.
DPO estimates implicit rewards using the log-probability ratio between a policy model’s response and that of a supervised fine-tuned (SFT) model, enabling preference learning without an explicit reward function. However, this implicit reward may misalign with the log-probability metric during inference. Moreover, DPO’s reliance on both policy and SFT models significantly increases GPU usage, especially for LLMs. The DPO loss, derived from the Bradley-Terry model, can create training imbalances, as it does not ensure an increase in the probability of positive samples—potentially reducing both positive and negative probability simultaneously. Unlike IPO   Azar et al. (2024), which constrains probability variation but weakens response distinction, DPO also exhibits length bias, favoring longer responses due to preference label distribution inconsistencies Lu et al. (2024). This issue, common in multi-stage RLHF methods, allows models to exploit verbosity for higher rewards without improving output quality, often generating responses nearly twice as long as labeled data.
@@
  • Competitive performance: Despite being reference-free, our method demonstrates competitive performance when compared to DPO and its variants Hong et al. (2024a); Ethayarajh et al. (2024). This performance advantage is consistent across a variety of training setups and comprehensive instruction-following benchmarks, including AlpacaEval 2 Li et al. (2023) and Arena-Hard v0.1 Li et al. (2024).
@@ 2 Related Work
Reinforcement learning with human feedback (RLHF) often utilizes the Bradley-Terry model Bradley and Terry (1952) to estimate the probability of success in pairwise comparisons between two independently evaluated instances. Additionally, a reward model is trained to assign scores to these instances. Reinforcement learning algorithms, such as proximal policy optimization (PPO) Schulman et al. (2017), are used to train models to maximize the reward model’s score for the selected response, ultimately enabling LLMs to align with human preferences Stiennon et al. (2020); Ziegler et al. (2019). A notable example is InstructGPT Ouyang et al. (2022), which showcased the scalability and adaptability of RLHF in training instruction-following language models. Alternative approaches, such as reinforcement learning with language model feedback (RLAIF Lee et al. (2023)), may also serve as feasible substitutes for human feedback Bai et al. (2022); Sun et al. (2023). Nevertheless, RLHF encounters challenges, including the need for extensive hyperparameter tuning due to the instability of PPO Rafailov et al. (2024) and the sensitivity of the reward models Wang et al. (2024a). Consequently, there is a pressing demand for more stable preference alignment algorithms.
@@
Several techniques for preference alignment reduce the reliance on reinforcement learning. Direct Policy Optimization (DPO) Rafailov et al. (2024) is a method that integrates reward modeling with preference learning. And Identity Preference Optimization (IPO) Azar et al. (2024) is introduced to mitigate potential overfitting issues in DPO. In contrast to RLHF and DPO, an alternative approach called Kahneman-Tversky Optimization (KTO) Ethayarajh et al. (2024) is proposed, which does not require pairwise preference datasets. Additionally, Preference Ranking Optimization (PRO) Song et al. (2024) introduces the incorporation of the softmax values from the reference response set into the negative log-probability (NLL) loss, allowing for a unified approach to supervised fine-tuning and preference alignment.
@@
Due to the reliance of DPO and DPO-like methods on both the policy model and the SFT model during the alignment process, they impose greater demands on GPU resources. Several techniques have been developed to alleviate this GPU requirement by eliminating the need for a reference model. CPO Xu et al. (2024) demonstrates that the ideal loss function without a reference model can serve as the upper bound of the DPO loss, with the SFT loss acting as a replacement for the KL divergence. ORPO Hong et al. (2024a) models the optimal reward as a log-odds function, removing the need for an additional fixed reference model. MaPO Hong et al. (2024b) builds on the ORPO approach by introducing a margin-aware term for aligning diffusion models without a reference model. SimPO Meng et al. (2024) adopts a similar reference-free preference learning framework as CPO but with improved stability due to its specific length normalization and target reward margin, leading to superior performance in various benchmarks.
@@ 3.1 Direct Preference Optimization (DPO)
We derive our method by first revisiting Direct Preference Optimization (DPO), which offers a simplified optimization objective within the RLHF framework. DPO operates on a dataset $\mathcal{D}={(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})}{i=1}^{N}$, where each input $x$ is paired with a preferred output $y_{w}$ and a less preferred one $y_{l}$. The loss function is defined as a form of maximum likelihood estimation for a policy $\pi\theta$:
@@
Here, $\pi_{\text{ref}}$ denotes a reference policy (typically from SFT), $\sigma$ is the sigmoid function, and $\beta$ is a scaling factor. This loss stems from a reparameterization of the reward and optimal policy, inspired by PPO. Unlike PPO, however, DPO enables training via supervised learning using only static preference-labeled data, avoiding the need for online environment interaction.
3.2 Improvement of DPO
Bradley-Terry model with home-field advantage
In Section 3.1, DPO employs the Bradley-Terry model, a well-established statistical framework frequently utilized in the analysis of competitive events, such as sporting contests. The Bradley-Terry model formalizes the human preference distribution $p^{*}$ as:
@@ 3.2 Improvement of DPO
The Bradley-Terry (BT) model utilized in DPO adopts its original formulation. However, several variants have been proposed to enhance the model’s capabilities. Notably, the Rao-Kupper model extends the BT framework by accounting for tied preferences, modeling the probability $p^{*}(y_{w}=y_{l}\mid x)$, which signifies that two responses, $(y_{w},y_{l})$, are deemed equivalent with respect to the given prompt $x$.
To better differentiate between the two responses, we reinterpret the loss response within the BT model as the "home team" in a competitive setting. Furthermore, to incorporate a potential home-field advantage, we introduce an intercept term $h$, refining the model’s capacity to capture systematic biases:
@@
Removal of reference model
For DPO, $d(x,y_{w},y_{l})$ represents the term within the function $\sigma$, as outlined in Section 3.1. DPO has been widely adopted in modern models. However, despite its advantages, DPO exhibits significant drawbacks compared to standard supervised fine-tuning, such as more memory consumption and substantial computational inefficiencies due to the usage of the reference model. These limitations underscore the critical need for exploring reference model-free RLHF approaches.
A recent approach, SimPO, utilizes the average log-likelihood function as a substitute for the reference model. However, the rationale behind this substitution remains insufficiently explained. In this work, we provide a detailed explanation to address this gap.
A recent method, CPO, demonstrates that when the reference policy $\pi_{\text{ref}}$ is set to $\pi_{w}$—an ideal policy that perfectly aligns with the true data distribution of preferred samples—the DPO loss $\mathcal{L}(\pi_{\theta};\pi_{w})+C$ is upper-bounded by $\mathcal{L}(\pi_{\theta};U)$, where $C$ is a constant. Building on this result, we approximate $d(x,y_{w},y_{l})$ using a uniform reference model:
@@
Next, in DPO and CPO, the implicit reward is defined as the log ratio of the response probability between the current policy model and the SFT model. However, this reward formulation does not directly align with the metric guiding generation, which is roughly the average log probability of a response generated by the policy model. This discrepancy between the training and inference phases may negatively impact performance. To address this, we replace the log probability with the average log probability in Eq. 4:
@@ 3.3 Length-Controlled Margin-Based Loss
Here, $m(y_{w},y_{l},x)$ denotes the margin quantifying the preference strength between the winning response $y_{w}$ and the losing response $y_{l}$ given input $x$, and $\lambda$ is a scaling factor. The function $r^{*}(x,y)$ provides the reward score for response $y$ conditioned on input $x$. Incorporating this margin enables the model to better differentiate reward scores, especially when the quality gap between responses is large.
Recent works have adopted this formulation to improve performance. For instance, the reward models in Llama-2-Chat (Touvron et al., 2023) and UltraRM (Cui et al., 2023) use discrete preference scores as margin terms, while SimPO (Meng et al., 2024) employs a fixed margin to ensure the preferred response always receives a higher reward than the less favored one. Nevertheless, issues such as length bias remain.
To address these challenges, we propose the Length-Controlled Margin-Based Loss. This loss explicitly regulates the length of generated responses, mitigating the tendency of large language models to prefer longer outputs. It also controls the probability decrease for both selected and rejected responses, enhancing the model’s ability to distinguish between correct and incorrect answers. Importantly, it enlarges the margin between the probabilities of chosen and rejected responses, strengthening the model’s discrimination of response quality. The full formulation is given below.
@@
In neural machine translation (NMT) adequacy evaluation, the use of the "power of 5" in the margin term has been shown to be the most effective approach. Prior studies Miao et al. (2021) have demonstrated its superiority through ablation experiments comparing various margin formulations. Additionally, the "power of 5" margin has been incorporated into recent Mixture-of-Experts (MoE) models, such as MoE-Summ Chen et al. (2024), achieving significant improvements across multiple tasks. Motivated by these findings, we adopt the "power of 5" margin term in this work.
@@
Objective. Finally, we obtain the LMPO final loss function by incorporating the above considerations:
@@
In summary, LMPO employs an implicit reward formulation that directly aligns with the generation metric, eliminating the need for a reference model. Next, it introduces a margin term $m(\mathbf{x},\mathbf{y}^{w},\mathbf{y}^{l})$ with two kinds of normalization methods to help separate the winning and losing responses, alleviate length bias and wining response probability decrement problems. Details of Z-score normalization and further analysis of LMPO loss are shown in Appendix A.
@@ 4.1 Experimental Setup
Models and Training Settings. We optimize preferences using two model families: Llama3-8B (AI@Meta, 2024) and Mistral-7B (Jiang et al., 2023), under two setups: Base and Instruct.
In the Base setup, following SimPO, we use pre-trained SFT models Zephyr-7B-SFT (Tunstall et al., 2023) and Llama-3-Base-8B-SFT as initialization. Preference optimization is then performed on the UltraFeedback dataset (Cui et al., 2023), which contains feedback from LLMs of varying quality.
For the Instruct setup, we use instruction-tuned models Mistral-7B-Instruct-v0.2 and Meta-Llama-3-8B-Instruct as SFT models. We adopt the same training data as SimPO: princeton-nlp/llama3-ultrafeedback and princeton-nlp/mistral-instruct-ultrafeedback for Llama3-8B and Mistral-7B, respectively.
These settings reflect recent advances, placing our models among top performers on several leaderboards.
@@
Additionally, we evaluate the models on ten downstream tasks in the Huggingface Open Leaderboard V1 and V2, following SimPO (Meng et al., 2024) and SIMPER Xiao et al. (2025). These downstream tasks include the AI2 Reasoning Challenge (25-shot) (Clark et al., 2018), TruthfulQA (0-shot) (Lin et al., 2021), Winogrande (5-shot) (Sakaguchi et al., 2021), GSM8K (5-shot) (Cobbe et al., 2021), IFEval Zhou et al. (2023), BBH Suzgun et al. (2022), MATH Hendrycks et al. (2021), GPQA Rein et al. (2024), MuSR Sprague et al. (2023), MMLU-PRO Wang et al. (2024b). We report the match accuracy for these conditional benchmarks. Additional details are provided in Appendix B.
Baselines We perform a comparative analysis of our method against several state-of-the-art offline preference optimization techniques, including DPO (Rafailov et al., 2024), SLiC Zhao et al. (2023), IPO (Azar et al., 2024), CPO (Xu et al., 2024), KTO (Ethayarajh et al., 2024) and SimPO (Meng et al., 2024). For SimPO, we use the model provided for the Llama3-8B family and replicate the SimPO methodology for the Mistral-7B family in our environment. For the other methods, we report the results provided by SimPO. We also tune the hyperparameters for SimPO and report the best performance achieved.